# EvalAlign: Supervised Fine-Tuning Multimodal LLMs with Human-Aligned Data for Evaluating Text-to-Image Models

### v2

@@ 3.3 Dataset Statistics
To summarize, we generate 24k images from 3k prompts based on 8 text-to-image models, which includes DeepFloyd IF [1], SD15 [40], LCM [28], SD21 [40], SDXL [33], Wuerstchen [32], Pixart [6], and SDXL-Turbo [48]. After data filtering, 4.5k images were selected as annotation data for task of text-image alignment. Subsequently, these images were carefully annotated to generate 13.5k text-image pairs, where 11.4k were used to the training dataset and 2.1k to the validation dataset. For the image faithfulness task, we select 12k images for annotate, producing 36k text-image pairs, with 30k are used to the training dataset and 6.2k to the validation dataset. Additionally, we employed 24 text-to-image models to generate 2.4k images from 100 prompts. After annotation, these images were used as testing dataset. Figure 3 and Figure 3 show the distribution of objects in different categories within our prompts, demonstrating the diversity and balance of our prompts.
@@ 4 Training and Evaluation Methods
4.1 Supervised Fine-Tuning the MLLM
As we mentioned above, we use MLLMs as the evaluation models and let it to answer a set of carefully-designed instructions, thereby achieving quantitative measurement of fine-grained text-to-image generation skills. Due to training bias, zero-shot MLLMs perform poorly when it comes to fine-grained evaluation on synthesized images, particularly in term of image faithfulness. To address, we conduct SFT training on the annotated data to align the MLLM with human preference prior on fine-grained image assessing. Formally, the SFT training sample can be denoted as a triplet, i.e., , where denotes the question (or the instruction), the multimodal input and the answer, respectively. Specifically, is a question about a fine-grained aspect of the generated image as well as the available answer options, is mainly the image and necessary textual description, while is the specific answer chosen from the given options and annotated by human. Notably, we expand each given option to a more detailed and descriptive version, benefiting the SFT performance by allowing the MLLM to understand the meaning of each option. During SFT training, the optimization objective is the original autoregressive loss function utilized to train LLMs, but calculated only on . Specifically, the loss function can be formulated as follows, where is the length of the ground truth answer:
@@ 4.2 Evaluation and Metrics
During inference, the MLLM, parameterized by , is required to generate the appropriate response , given the specific question and the multimodal input in an autoregressive way:
This autoregressive generation procedure is considered complete once the model generates a EOS token or the generated response exceeds the preset maximum generation length. After the generation, we use rule-based filtering and regular expression to extract the option chosen by the MLLM, where each option has an exclusive predefined score to quantitatively measure a fine-grained skill specified by the question :
where represents the procedure of option extraction and score searching. We devise two holistic and detailed question sets, and , that encompass every aspect of image faithfulness and text-image alignment, respectively. Consequently, Our metric EvalAlign can be defined by simply summing the scores of the two question sets:
where and indicates the image faithfulness score and image-text alignment score evaluated by EvalAlign.
@@ 4.3 Implementation Details
We leverage the LLaVA-NeXT model and applied the LoRA fine-tuning method to align it with the training dataset as described previously. The entire training process was conducted on NVIDIA A100 GPUs with hours, using a learning rate of 510-5. Additionally, we restricted the LoRA fine-tuning to the Q and K weights of the attention module, as extending the fine-tuning to the ViT and projection modules was found to cause overfitting. Importantly, our method does not necessitate any modifications to the MLLM, rendering it compatible with any MLLM model. For the ablation study, we evaluated the LLaVA-NeXT 13B model on the validation dataset. In the final experiment, we apply supervised fine-tuning to the LLaVA-NeXT 34B model on the testing dataset and compare against different evaluation metrics.
@@ 5 Experimental Results
In this section, we show results of EvalAlign on benchmarking 24 text-to-image generative models. We also compare with popular text-to-image metrics and frameworks, including Inception Score (IS) [44], Frechet Inception Distance (FID) [15], CLIP-score [14], and HPS v2 [54]. Please refer to supplementary materials for detailed results and analyses.
@@ 5.1 Main Results
Model | Human | EvalAlign |
PixArt XL2 1024 MS [6] | 2.2848 | 1.6415 |
Dreamlike Photoreal v2.0 [11] | 2.0070 | 1.4522 |
SDXL Refiner v1.0 [47] | 1.9229 | 1.6072 |
SDXL v1.0 [33] | 1.8136 | 1.4675 |
Wuerstchen [32] | 1.7837 | 1.4279 |
LCM SDXL [28] | 1.6910 | 1.3391 |
Openjourney [34] | 1.6667 | 1.175010 |
Safe SD MAX [31] | 1.6491 | 1.2175 |
LCM LORA SDXL [28] | 1.6387 | 1.3833 |
Safe SD STRONG [31] | 1.630810 | 1.146611 |
Safe SD MEDIUM [31] | 1.627511 | 1.129815 |
Safe SD WEAK [31] | 1.607812 | 1.118817 |
SD v2.1 [40] | 1.552413 | 1.109418 |
SD v2.0 [40] | 1.527714 | 1.130014 |
Openjourney v2 [35] | 1.500015 | 0.995620 |
Redshift diffusion [39] | 1.473316 | 1.138212 |
Dreamlike Diffusion v1.0 [10] | 1.465217 | 1.2052 |
SD v1.5 [40] | 1.441718 | 1.136213 |
IF-I-XL v1.0 [1] | 1.380819 | 0.922122 |
SD v1.4 [40] | 1.359220 | 0.951121 |
Vintedois Diffusion v0.1 [51] | 1.356221 | 1.079719 |
IF-I-L v1.0 [1] | 1.263522 | 0.881423 |
MultiFusion [29] | 1.237223 | 1.129816 |
IF-I-M v1.0 [1] | 1.013524 | 0.792824 |
Model | Human | EvalAlign |
IF-I-XL v1.0 [1] | 5.4500 | 5.5300 |
IF-I-L v1.0 [1] | 5.2300 | 5.4500 |
SDXL Refiner v1.0 [47] | 5.2100 | 5.4000 |
LCM SDXL [28] | 5.1800 | 5.3300 |
PixArt XL2 1024 MS [6] | 5.1100 | 5.3100 |
IF-I-M v1.0 [1] | 5.0800 | 5.2200 |
LCM LORA SDXL [28] | 5.0600 | 5.2700 |
SDXL v1.0 [33] | 5.0300 | 5.3500 |
Wuerstchen [32] | 4.8700 | 5.1700 |
Openjourney [34] | 4.830010 | 4.920015 |
SD v2.1 [40] | 4.800011 | 5.070011 |
MultiFusion [29] | 4.680012 | 4.800018 |
Dreamlike Diffusion v1.0 [10] | 4.660013 | 5.150010 |
SD v2.0 [40] | 4.640014 | 5.010012 |
Vintedois Diffusion v0.1 [51] | 4.620015 | 4.950014 |
Safe SD STRONG [31] | 4.600016 | 4.830017 |
Dreamlike Photoreal v2.0 [11] | 4.560017 | 4.980013 |
Safe SD WEAK [31] | 4.530018 | 4.710020 |
SD v1.4 [40] | 4.520019 | 4.760019 |
SD v1.5 [40] | 4.450020 | 4.900016 |
Safe SD MEDIUM [31] | 4.400021 | 4.560024 |
Redshift diffusion [39] | 4.350022 | 4.670021 |
Safe SD MAX [31] | 4.310023 | 4.590023 |
Openjourney v2 [35] | 4.150024 | 4.650022 |
@@
Evaluation on image faithfullness. We conducted the image faithfullness evaluation on the testing dataset, as detailed in Table 3. Our evaluation metric, EvalAlign, demonstrates close alignment with human preferences Specifically, the rankings of the top and bottom 10 models by both EvalAlign and human evaluation scores show remarkable consistency, confirming that our metric closely mirrors human evaluation. Importantly, most of the images from the 24 generative models were not present during training, showcasing the robust generalization capability of our evaluation method.
Evaluation on text-image alignment. The evaluation of text-image alignment on the testing dataset was conducted follow the same manner of image faithfullness evaluation. Table 3 reveals that the rankings based on the EvalAlign metric and human evaluation metric for the 24 models are generally consistent. This consistency likely stems from the proximity of MLLM’s pre-training tasks to the text-image alignment evaluation tasks, resulting in better performance on text-image alignment evaluation task.
@@ 5.2 Ablations and Analyses of EvalAlign
Results on different prompt categories. SFT significantly enhances the performance of MLLMs on evaluation tasks. We conducted experiments comparing the LLava-Next 13B model with and without SFT. As shown in Table 5 and Table 5, the results demonstrate that SFT considerably improves performance across all prompt categories in both image faithfulness and text-to-image alignment, closely aligning the MLLM’s predictions with human evaluations. Note that Table 5 shows that the baseline method without SFT performs poorly on the image faithfulness task, which suggests the necessity of SFT.
@@
Effect of dataset size for vision-language model training. In order to explore the effects of data size, we train model on image faithfulness evaluation task under three different data size: 200, 500 and 800. As illustrated in Table 6, performance generally enhances with more training data. Notably, training with just 500 data nearly maximizes accuracy, with further increases to 800 data yielding only slight improvements. This result suggests that our method requires only a small amount of annotated data to achieve good performance, highlighting the cost-effectiveness of our approach.
Effect of model size. Transformers have demonstrated a strong capacity to scale effectively across a variety of tasks, as noted in key studies [9, 36]. In Table 7, we illustrate the advantages of scaling up the MLLM for image faithfulness evaluation. Specifically, increasing the MLLM model size from 7B to 34B results in notable improvements in performance.

### v3

@@ 3.3 Dataset Statistics
To summarize, we generate 24k images from 3k prompts based on 8 text-to-image models, which includes DeepFloyd IF (Alex Shonenkov & et al., 2023), SD15 (Rombach et al., 2022a), LCM (Luo et al., 2023), SD21 (Rombach et al., 2022a), SDXL (Podell et al., 2023), Wuerstchen (Pernias et al., 2023), Pixart (Chen et al., 2023b), and SDXL-Turbo (Stability AI, a). After data filtering, 4.5k images are selected as annotation data for task of text-image alignment. Subsequently, these images are carefully annotated to generate 13.5k text-image pairs, where 11.4k are used to the training dataset and 2.1k to the validation dataset. For the image faithfulness task, we select 12k images for annotation, yielding 36k text-image pairs, with 30k are used to the training dataset and 6.2k to the validation dataset. Additionally, we employed 24 text-to-image models to generate 2.4k images from 100 prompts. After annotation, these images are used as testing dataset. Figure 3 and Figure 3 show the distribution of objects in different categories within our prompts, demonstrating the diversity and balance of our prompts.
@@ 4 Training and Evaluation Methods
4.1 Supervised Finetuning the MLLM
As we mentioned above, we use MLLMs as the evaluation models and let it to answer a set of carefully-designed instructions, thereby achieving quantitative measurement of fine-grained text-to-image generation skills. Due to data bias, zero-shot MLLMs perform poorly when it comes to evaluation on generated images, particularly in term of image faithfulness. To solve this problem, we apply SFT training on the detailed human annotation to align the MLLM with human preference prior. Formally, the SFT training sample can be denoted as a triplet: question (or the instruction), multimodal input and answer. During SFT training, the optimization objective is the autoregressive loss function utilized to train LLMs, but calculated only on the answer, the loss function can be formulated as follows:
where is the length of the ground truth answer, is a fine-grained question of the generated image and its available answer, is the image and textual description, while is the human annotated answer selected from the given options. Notably, we expand each option to make it more detailed and descriptive, thereby benefiting SFT performance by allowing the MLLM to better understand the meaning of each option.
@@ 4.2 Evaluation and Metrics
To evaluate synthesized images with consideration of its synthetic nature, EvalAlign is designed to evaluate image faithfulness and text-image alignment in a fine-grained way. Notably, image faithfulness and text-image alignment are two common errors occurred in synthesized images, whereas real images inherently exhibit high levels of both image faithfulness and text-image alignment.
Image Faithfulness measures whether synthesized images are faithful to real-world commonsense. With higher image faithfulness, the visual elements of generated images more closely resemble their real-world counterparts. Unfortunately, text-to-image models often generate images with low faithfulness, such as distorted body structures and human hands. This is also a critical reason why we set image faithfulness as one of the benchmarking aspects in EvalAlign. Additionally, evaluating image faithfulness requires considering the input prompts, as prompts may describe unreal or impossible scenarios that inherently affect the faithfulness of the generated images. For example, when prompts like "a dog walking like a human" or "a man on Mars without a spacesuit" are provided, the generated images may naturally deviate from real-world image faithfulness. Under such circumstances, the synthesized images cannot be regarded as low in image faithfulness since the generative models are merely following prompts that contain super-reality scenarios.
Text-Image Alignment evaluates whether generated images are aligned with their conditioned prompts. In the inference settings of text-to-image models, the image generation process is conditioned on textual prompts, requiring alignment between the text prompts and the synthesized images. However, through our observations, text-to-image models cannot consistently follow input prompts, often yielding images with visual elements misaligned with the input prompts. For example, models may generate images featuring an orange cat when conditioned on the text prompt "a blue cat."
During inference, the multimodal large language model (MLLM) is required to generate an appropriate response given a specific question and multimodal input in an autoregressive manner:
where is the -th generated token, represents the sequence of tokens generated before step , and denotes the parameters of the fine-tuned MLLM. This autoregressive generation process is considered complete once the model generates an end-of-sequence (EOS) token or the generated response exceeds a preset maximum generation length. After generation, we employ rule-based filtering and regular expressions to extract the option chosen by the MLLM. Each option is assigned a unique predefined score to quantitatively measure a fine-grained skill specified by the question :
where represents the procedure of option extraction and score mapping.
We devise two holistic and detailed question sets, and , that encompass every aspect of image faithfulness and text-image alignment, respectively. Consequently, our metric, EvalAlign, can be defined by averaging the scores of the questions in the two sets:
where and indicate the image faithfulness score and the text-image alignment score evaluated by our method, respectively.
@@ 4.3 Implementation Details
For details about the SFT training, we apply LoRA (Hu et al., 2021) finetuning on LLaVA-NeXT (Liu et al., 2024a) models to align them with the EvalAlign dataset. Additionally, we merely adapt LoRA finetuning on the Q and K weights of the attention module, as extending the finetuning to the ViT (Dosovitskiy, 2020) and projection modules will lead to overfitting. The entire training process is conducted on NVIDIA A100 GPUs for hours, with a learning rate of 510-5. As for the ablation study, we evaluate the finetuned LLaVA-NeXT 13B model on the validation dataset. In the final experiment, we apply SFT to the LLaVA-NeXT 34B model on the testing dataset to testify its generalization ability.
@@ 5.1 Main Results
Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore |
PixArt XL2 1024 MS (Chen et al., 2023b) | 2.2848 | 1.6415 | 31.6226 | 0.8580 | 0.9696 | 22.1335 |
Dreamlike Photoreal v2.0 (dreamlike.art, b) | 2.0070 | 1.4522 | 29.2322 | 0.828612 | 0.1886 13 | 21.2271 |
SDXL Refiner v1.0 (Stability AI, b) | 1.9229 | 1.6072 | 29.8197 | 0.8566 | 0.7245 | 22.0492 |
SDXL v1.0 (Podell et al., 2023) | 1.8136 | 1.4675 | 29.0620 | 0.8467 | 0.7043 | 21.8106 |
Wuerstchen (Pernias et al., 2023) | 1.7837 | 1.4279 | 30.6622 | 0.819914 | 0.3212 11 | 21.3720 |
LCM SDXL (Luo et al., 2023) | 1.6910 | 1.3391 | 29.3588 | 0.833510 | 0.5304 | 21.6532 |
Openjourney (PromptHero, a) | 1.6667 | 1.175010 | 26.347513 | 0.819615 | 0.1478 16 | 20.863710 |
Safe SD MAX (Patrick et al., 2022) | 1.6491 | 1.2175 | 25.739617 | 0.755524 | -0.050722 | 20.459421 |
LCM LORA SDXL (Luo et al., 2023) | 1.6387 | 1.3833 | 27.329910 | 0.8364 | 0.4959 | 21.4824 |
Safe SD STRONG (Patrick et al., 2022) | 1.630810 | 1.146611 | 25.576418 | 0.816518 | -0.102223 | 20.621118 |
Safe SD MEDIUM (Patrick et al., 2022) | 1.627511 | 1.129815 | 26.279814 | 0.810120 | 0.2042 12 | 20.788012 |
Safe SD WEAK (Patrick et al., 2022) | 1.607812 | 1.118817 | 26.118015 | 0.780923 | -0.126424 | 20.387324 |
SD v2.1 (Rombach et al., 2022a) | 1.552413 | 1.109418 | 26.582312 | 0.8377 | 0.4116 | 21.0502 |
SD v2.0 (Rombach et al., 2022a) | 1.527714 | 1.130014 | 25.348121 | 0.817017 | 0.0872 18 | 20.752913 |
Openjourney v2 (PromptHero, b) | 1.500015 | 0.995620 | 24.698423 | 0.795822 | -0.041521 | 20.408822 |
Redshift diffusion (Redshift-Diffusion, ) | 1.473316 | 1.138212 | 25.157222 | 0.810121 | 0.0218 20 | 20.615519 |
Dreamlike Diffusion v1.0 (dreamlike.art, a) | 1.465217 | 1.2052 | 29.6506 | 0.8543 | 0.6508 | 21.2664 |
SD v1.5 (Rombach et al., 2022a) | 1.441718 | 1.136213 | 25.497219 | 0.821413 | 0.1686 14 | 20.714316 |
IF-I-XL v1.0 (Alex Shonenkov & et al., 2023) | 1.380819 | 0.922122 | 27.4512 | 0.8449 | 0.6087 | 20.747414 |
SD v1.4 (Rombach et al., 2022a) | 1.359220 | 0.951121 | 25.369720 | 0.819016 | 0.1050 17 | 20.653517 |
Vintedois Diffusion v0.1 (Vintedois-Diffusion v0., 1) | 1.356221 | 1.079719 | 26.590111 | 0.8341 | 0.3562 10 | 20.835811 |
IF-I-L v1.0 (Alex Shonenkov & et al., 2023) | 1.263522 | 0.881423 | 27.4836 | 0.8384 | 0.4463 | 20.717015 |
MultiFusion (Marco et al., 2023) | 1.237223 | 1.129816 | 23.813324 | 0.815119 | 0.0695 19 | 20.478020 |
IF-I-M v1.0 (Alex Shonenkov & et al., 2023) | 1.013524 | 0.792824 | 25.952216 | 0.832911 | 0.1637 15 | 20.403523 |
Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore |
IF-I-XL v1.0 (Alex Shonenkov & et al., 2023) | 5.4500 | 5.5300 | 32.547710 | 0.8579 | 0.4391 | 21.199810 |
IF-I-L v1.0 (Alex Shonenkov & et al., 2023) | 5.2300 | 5.4500 | 32.7140 | 0.8538 | 0.3820 | 21.128412 |
SDXL Refiner v1.0 (Stability AI, b) | 5.2100 | 5.4000 | 35.6465 | 0.8528 | 0.4738 | 22.3532 |
LCM SDXL (Luo et al., 2023) | 5.1800 | 5.3300 | 33.8011 | 0.8512 | 0.3833 | 21.9620 |
PixArt XL2 1024 MS (Chen et al., 2023b) | 5.1100 | 5.3100 | 37.0493 | 0.8634 | 0.6542 | 22.3926 |
IF-I-M v1.0 (Alex Shonenkov & et al., 2023) | 5.0800 | 5.2200 | 31.095114 | 0.8434 | 0.0499 10 | 20.827020 |
LCM LORA SDXL (Luo et al., 2023) | 5.0600 | 5.2700 | 32.7752 | 0.834910 | 0.1618 | 21.7627 |
SDXL v1.0 (Podell et al., 2023) | 5.0300 | 5.3500 | 35.1593 | 0.8540 | 0.4322 | 22.1291 |
Wuerstchen (Pernias et al., 2023) | 4.8700 | 5.1700 | 36.4632 | 0.8381 | 0.2513 | 21.7779 |
Openjourney (PromptHero, a) | 4.830010 | 4.920015 | 31.149512 | 0.817316 | -0.086714 | 21.116313 |
SD v2.1 (Rombach et al., 2022a) | 4.800011 | 5.070011 | 31.101713 | 0.827814 | -0.045312 | 21.2093 |
MultiFusion (Marco et al., 2023) | 4.680012 | 4.800018 | 28.795724 | 0.826415 | -0.133715 | 20.962517 |
Dreamlike Diffusion v1.0 (dreamlike.art, a) | 4.660013 | 5.150010 | 34.8196 | 0.8493 | 0.2295 | 21.5550 |
SD v2.0 (Rombach et al., 2022a) | 4.640014 | 5.010012 | 30.615317 | 0.829813 | -0.142416 | 21.190511 |
Vintedois Diffusion v0.1 (Vintedois-Diffusion v0., 1) | 4.620015 | 4.950014 | 31.950311 | 0.831912 | -0.022211 | 21.114114 |
Safe SD STRONG (Patrick et al., 2022) | 4.600016 | 4.830017 | 30.661516 | 0.775123 | -0.502822 | 20.749121 |
Dreamlike Photoreal v2.0 (dreamlike.art, b) | 4.560017 | 4.980013 | 33.7712 | 0.834411 | -0.085913 | 21.4832 |
Safe SD WEAK (Patrick et al., 2022) | 4.530018 | 4.710020 | 30.564418 | 0.814018 | -0.272818 | 20.989916 |
SD v1.4 (Rombach et al., 2022a) | 4.520019 | 4.760019 | 29.914920 | 0.804820 | -0.343819 | 20.846219 |
SD v1.5 (Rombach et al., 2022a) | 4.450020 | 4.900016 | 30.167319 | 0.814217 | -0.221317 | 20.864018 |
Safe SD MEDIUM (Patrick et al., 2022) | 4.400021 | 4.560024 | 30.782015 | 0.797421 | -0.359120 | 21.025715 |
Redshift diffusion (Redshift-Diffusion, ) | 4.350022 | 4.670021 | 29.286522 | 0.806619 | -0.417221 | 20.632723 |
Safe SD MAX (Patrick et al., 2022) | 4.310023 | 4.590023 | 29.812621 | 0.760124 | -0.609524 | 20.704622 |
Openjourney v2 (PromptHero, b) | 4.150024 | 4.650022 | 29.238923 | 0.785122 | -0.605123 | 20.597324 |
@@
Evaluation on image faithfulness. We evaluate image faithfulness on the testing dataset to ensure that the finetuned MLLM aligns with human judgment and generalizes to unseen data. As detailed in Table 2, the finetuned MLLM successfully aligns with human preferences on image faithfulness, indicating its ability of image faithfulness evaluation is close to human. Specifically, the rankings of the top and bottom 10 models by both EvalAlign and human evaluation scores are remarkably consistent. Besides, most of the images in the testing dataset, especially those from the 16 exclusive generative models, are not present during the SFT training, showcasing the robust generalization capability of our models.
Evaluation on text-image alignment. The evaluation of text-image alignment on the testing dataset is similar to that of image faithfulness. Table 2 reveals that the rankings of the 24 evaluated models by EvalAlign are generally consistent with human annotators. We believe that the consistency on image faithfulness and text-image alignment evaluations mainly stems from our annotated high-quality SFT dataset. It also proves that, with the annotated dataset and the extraordinary image-text joint understanding ability owned by MLLMs, we can easily finetune a MLLM to conduct the evaluation with low cost but close-to-human performance.
@@ 5.2 Ablations and Analyses of EvalAlign
Results on different prompt categories. Since MLLMs are not specifically trained to perform evaluations, they are naturally ill-suited for this task, hindering their task performances. Therefore, we need to annotate SFT data for this task and finetune the MLLMs accordingly. To verify the necessity, We conduct experiments comparing the LLava-Next 13B model with and without SFT. As shown in Table 5 and Table 5, the results demonstrate that SFT training considerably improves performance across all prompt categories in both image faithfulness and text-to-image alignment, closely aligning the MLLM’s predictions with human evaluations. Note that Table 5 illustrates that the baseline method without SFT performs poorly in image faithfulness and text-image alignment evaluations, particularly in the former.
@@
Effect of training dataset size for vision-language model training. In order to explore the effects of data size and determine the sufficient amount of training data, we train the model on image faithfulness evaluation task with images and their annotations sourced from 200, 500 and 800 prompts. As illustrated in Table 6, the evaluation performance continuously enhances as more training data is used. Notably, training with just 500 prompts nearly maximizes accuracy, with further increases to 800 data yielding only marginal improvements. This result suggests that our method requires only a small amount of annotated data to achieve good performance, highlighting its cost-effectiveness. Generally, since more data leads to better performance, we use all of the available data to finetune our models and release this data to the research community to bootstrap further study.
Effect of model size. Since transformers are known for their scalability (Radford et al., 2018; Dehghani et al., 2023), we investigate the effect of the model size on the performance of image faithfulness evaluation. As illustrated in Table 7, the benefits of scaling up the utilized MLLMs are remarkably significant, where increasing the model size from 7B to 34B results in substantial improvements in evaluation performance. For this consequence, for the final version of the EvalAlign evaluation model, we choose LLaVA-NeXT 34B, the largest model in LLaVA-NExT series, and finetune it on our meticulously curated SFT data. Since some users of EvalAlign cannot afford MLLM inference with 34B parameters, we will make the 13B and 34B models publicly available.